Circular No. 5820

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

URANUS W. Wild, B. Carter, and E. Kibblewhite, University of Chicago; and D.

Wittman, M. Lloyd-Hart, and R. Angel, University of Arizona, report: “On May 30.476-30.493 UT, a series of J, H, and short K images (resolution ∼ 0′′.5) were taken of Uranus with the 1.8-m MMT E mirror, utilizing the University of Arizona ACME adaptive optics system. The J and H images were of 30-s, and the K of 60-s, duration. The adaptive optics system, in the configuration used, stabilizes the centroid of the image at a 50-Hz sampling rate using the signal in the visible portion of the spectrum. A dark region appears in all the compensated images that were acquired; we note the the images were taken on various portions of the NICMOS array. The dark spot appears to be larger in the J-band image. Additional features in the image are significant limb brightening in the polar region and what appears to be an irregular dark band near the pole. The NICMOS plate scale is 0′′.2/pixel and the dark spot subtends ∼ 2-3 pixels in the J band and 2 pixels in the H and K bands, at the limit of resolution. The spot is estimated to be at latitude −10^{o} on the planet (in the J band image). In our images, with appropriate stretching of the signal, which also brings out the sky noise, it is possible to discern the rings in each band. We judge the rings in the images to be real — as opposed to the seeing disk of the planet — because they are smaller east-west than north-south (i.e., their size and orientation appear to be correct). Unfortunately, we did not get additional imagery beyond this time interval, so we were unable to observe these features rotate on the planet. Further observations are encouraged to verify the presence and evolution of the dark spot on the planet.”

SUPERNOVA 1993J IN NGC 3031 J. M. Marcaide, Universitat de Valencia; A. E. E. Rogers, Haystack

Observatory; A. Alberdi, J. C. Guirado, and A. Rius, Consejo Superior de Investigaciones Cientificas, Spain; P. Elosegui, I. I. Shapiro, and M. I. Ratner, Harvard-Smithsonian Center for Astrophysics; E. Perez, Instituto de Astrofisica de Canarias; and A. R. Whitney, Haystack Observatory, report an improved position for SN 1993J from VLBI observations made during Apr. 24.90-25.06 UT: α = 9^{h}55^{m}24^{s}.7740 ± 0^{s}.0006, δ = +69^{o}01′13′′.700 ± 0′′.003 (equinox J2000.0)

1993 June 24 Daniel W. E. Green